Ford says that they expect that 20% of the Mustangs they sell to be Black , 30% Gray, & 50% Red. Conduct a x^2 test to determine if these observed frequencies match Ford's expectations at a 90% LC.

|    fa | fe

Black | 5 |    6

Gray |    9    |    9

Red |    16 | 15

Please show me how to get this answer...
Answer #1

Black = 5+6=11

Gray = 9 +9 = 18

Red = 16 +15 = 31

(5) Conclusion

It is concluded that the null hypothesis Ho is not rejected. Therefore, there is NOT enough evidence to claim that some of the population proportions differ from those stated in the null hypothesis, at the α=0.1 significance level.